# Present address: ]RIKEN Center for Quantum Computing (RQC), RIKEN, Wako-shi, Saitama 351-0198, Japan

### v1

@@ Abstract
We report the image current detection of electrons in a room-temperature Paul trap at microwave frequencies. By selectively leveraging distinct cavity modes for trapping and detection, our approach effectively extracts electron signals otherwise buried in the microwave drive used for pseudo-potential formation. When the trapped electrons resonate with the cavity mode, we observe a mode excitation and its exponential decay attributed to resistive cooling. Detuning electrons from the cavity resonance halts this decay, and sweeping electrons’ secular frequency reveals their oscillatory spectrum. Implementing this experiment at cryogenic temperatures could enable the image current detection and ground-state cooling of a single electron in Paul traps.
@@ I Introduction
Various trapping mechanisms have been explored to implement electron-based qubits in Penning traps. Planar Penning traps have gained attention due to their potential for two-dimensional scalability [26, 27, 28], which is required for quantum error correction algorithms such as surface codes [29]. However, their inherent lack of symmetry introduces significant anharmonicity in the trap potential, making image current detection of a single electron unfeasible [30]. While theoretical methods have been proposed to mitigate this issue [31], no successful single electron measurements have been reported.
@@
Unlike Penning traps, where an electron exhibits large-radius magnetron motion, Paul traps confine an electron closer to the trap center where anharmonicity is minimal. This localization enhances the feasibility of both single electron detection and scalable two-dimensional architectures, shifting the attention toward electron Paul traps [32, 33, 34, 35]. The Paul trapping of electrons was recently demonstrated [36], with detection achieved by ejecting electrons onto a micro-channel plate.
@@
Essentially, quantum information processing demands repeated operations, requiring the continuous confinement of electrons and their image current detection [5]. Equally crucial is ground-state cooling of the electron’s motion, a prerequisite for precise quantum manipulations [37]. However, both approaches present challenges: the former requires stringent filtering of the trap’s microwave field to extract the electron signal and protect detection circuits [35], while the latter demands resistive cooling at cryogenic temperatures, necessitating Paul traps inside a cryostat where cooling capacity is limited [36]. Overcoming these obstacles requires an integrated experimental framework combining advanced microwave filtering with power-efficient trap designs.
@@
Figure 1: Schematic illustration and characterization of the three-dimensional coupled coaxial cavities. (a) Design of the three-dimensional coupled coaxial cavities. (b) Cross-sectional view of the coupled coaxial cavities. The inset provides a magnified view of the pin geometry, where electrons are confined. (c) Radial ($r$-axis) electric field distribution of the out-of-phase $\lambda/4$ mode (readout mode) in cylindrical coordinates, used for microwave image current detection of trapped electrons. The inset highlights the strong electric field between the pins, forming a localized electric dipole field. (d) Radial electric field distribution of the in-phase $45\lambda/$ 4 mode (trapping mode), which generates the pseudo-potential. The inset illustrates the corresponding potential structure responsible for electron confinement. (e) Top: Pseudo-potential $U_{\mathrm{eff}}$ along the $z$-axis, centered at potential minimum (blue curve), compared to an ideal harmonic potential (dashed orange curve). Bottom: Deviation $\Delta U_{\mathrm{eff}}$ from the ideal harmonic potential, quantifying the anahronicity. The gray regions indicate the positions of the pin tips. (f) Pseudo-potential distribution and its deviations along the radial axis which reveal the influence of the cavity geometry on the potential shape.
@@
To address these challenges, we previously proposed three-dimensional coupled coaxial cavities [37]. This device can effectively extract electron signals from the trap’s microwave field by leveraging distinct resonance modes for pseudo-potential formation and electron detection. When integrated with superconducting materials, this cavity concentrates electric fields at a single point and has a very high $Q$ factor. Their synergy facilitates power-efficient electron Paul traps ($\sim\mathrm{mW}$) compatible with cryostat. In this work, we successfully realize electron Paul traps using coupled coaxial cavities. We achieve image current detection of electrons and observe exponential signal decay attributed to resistive cooling at room temperature. These results pave the way towards realizing image current detection and spin-readout of a single electron qubit at cryogenic temperature.
@@ II.1 Three-dimensional Coupled Coaxial Cavities
The pseudo-potential for the electron Paul trap is generated by oscillating quadrupole electric fields and can be formulated as $U_{\mathrm{eff}}=(q_{e}^{2}E^{2}/4m_{e}\Omega^{2})$. Here, $E$ and $\Omega$ represent the amplitude and angular frequency of the oscillating inhomogeneous electric field, while $q_{e}$ and $m_{e}$ are the charge and mass of the electron. Figure 1(e) and 1(f) show simulated pseudo-potentials along r- and z-axes when the pin voltage is $92\,\mathrm{V}$, which sets the electron oscillation to resonate with readout mode ($\omega_{z}=\omega_{1}^{\mathrm{out}}$). This pin voltage can be achieved with a microwave input power of $1.5\,\mathrm{W}$ under the current coupling condition and could be reduced to $800\,\mathrm{mW}$ if the critical coupling is achieved. The potential depth is approximately $0.8\,\mathrm{eV}$, limited by the confinement along the $z$-direction. The anharmonicity of the trap potential is illustrated as $\Delta_{\mathrm{eff}}$, the deviations from the ideal harmonic potentials (shown as orange dashed lines) in the bottom panels highlighted as green curves.
@@
We note that anharmonicity in the Pseudo-potential causes amplitude-dependent frequency shifts in the secular motion of trapped electrons. This effect broadens the motion spectrum and degrades the signal-to-noise ratio (SNR) in image current detection [30]. In our coupled coaxial cavities, anharmonicity primarily depends on the separation and the radius of the pin electrodes. For our experiment, we used a pin separation of $400\,\mathrm{\mu m}$ and a radius of $250\,\mathrm{\mu m}$. Considering the symmetry of the system, the pseudo-potential along the $z$-axis can be modeled using $U_{\mathrm{eff}}(z)=\frac{1}{2}m_{e}\omega_{z}^{2}(z^{2}+C_{4}z^{4}+C_{6}z^{6})$ ignoring odd-order terms. From the polynomial fit to the potential shown in Fig. 1(e), we obtained $C_{4}=-1.5\crossproduct 10^{-5}\,\mathrm{\mu m}^{-2}$ and $C_{6}\approx 0\,\mathrm{\mu m}^{-4}$. The amplitude-dependent secular frequency shifts are given by [31]:
@@ II.2 Electron Source
To effectively load electrons into traps, the kinetic energy of the electron source must be sufficiently lower than the trap depth, highlighting the importance of preparing a low-energy electron source. In the Penning trap system, thermionic emission or field emission points are known to be efficient for loading low-energy electrons into traps [40, 41]. In this method, high-energy primary electrons are generated and collide with the device wall, producing background gases. Those background gases collide with other high-energy primary electrons, producing lower-energy secondary electrons that are trapped efficiently.
This study used the photoelectron effect to generate high-energy primary electrons. Here, we directed a continuous multimode laser with a central wavelength of $405\,\mathrm{nm}$ and a power of $300\,\mathrm{mW}$ onto the pins inside the coupled coaxial cavities through a hole [see Fig. 2(a)] to produce primary photoelectrons. Although the energy of the laser at $405\,\mathrm{nm}$ (approximately $3.06\,\mathrm{eV}$) is theoretically insufficient to produce photoelectrons given the work function of gold plated on the device wall ($5.30\,\mathrm{eV}$) [42], biasing the pins with an oscillating electric field enables Schottky-enabled photoemission [43]. In addition, the multiple-photon process can also facilitate photoemission [44].
@@ II.3 Experimental Setup
The signal of trapped electrons was observed by measuring the excitation of the readout mode via coupling ports $P_{2}$ and $P_{3}$ using a spectrum analyzer. Here, filtering the trap MW field from the electron signals is crucial to protect the amplifier and the spectrum analyzer. Therefore, we implemented a three-stage filtering in the electronic circuit. Firstly, coupling ports $P_{2}$ and $P_{3}$ are positioned at the node of the trapping mode, resulting in a $30\,\mathrm{dB}$ reduction of trap MW leakage into the detection circuit [Fig. 1(d)]. Secondly, coupling ports $P_{2}$ and $P_{3}$ are connected to the Double Arrow $180^{\circ}$ Hybrid Coupler to induce destructive interference of the trap MW [Fig. 2(a)]. While the trapping mode excites these ports in phase, the readout mode excites them with a phase difference of $\pi$. Since the hybrid coupler processes the input signals by outputting their sum or difference, an appropriate choice of output ports can suppress the signal from the trapping mode by $16\,\mathrm{dB}$ while preserving the readout mode. Finally, the signal is transmitted through two low-pass filters with a cut-off frequency of $1.65\,\mathrm{GHz}$, eliminating the trap MW by $80\,\mathrm{dB}$ [Fig. 2(c)]. These result in a total filtering of approximately $126\,\mathrm{dB}$ for trap MW. After the filtering, the readout mode excitation signals are amplified by $62\,\mathrm{dB}$ with a low-noise amplifier followed by another amplifier before the signal analyzer [Fig. 2(c)].
@@ III.1 Image Current Detection
When the $V_{\mathrm{MW}}$ is appropriately tuned to resonate electrons’ center-of-mass (COM) motion with the readout mode, an excitation signal is observed. Figure 3(b) presents the results for sequence (i), where $V_{\mathrm{MW}}$ remains constant resonating electrons with the readout mode. The excitation signal appears and gradually decays with a time constant of $t_{\mathrm{cool}}=1.74\ \mathrm{s}$ before reaching a steady state. The observed excitation lasted over ten seconds even after electron loading finished, confirming successful electron trapping and suggesting successful image current detection.
@@ IV Conclusions and Outlook
The next challenge toward realizing electron qubits in Paul traps is the image current detection and spin-readout of a single electron [34]. However, the anharmonicity in the trap potential remains a major obstacle. In Penning traps, this issue is mitigated through tailored static electric field compensation [1]. In contrast, applying such compensation in Paul traps has been challenging due to the complexity of the nonlinear Mathieu equation and limited to systems with small stability parameters ($q\ll 0.1$) [47]. Recently, we proposed a novel approach that extends static electric field compensation to nonlinear Mathieu equations, enabling anharmonicity suppression of Paul traps even in high stability parameters ($q>0.5$) [48]. Implementing this method could improve electron confinement and facilitate single-electron detection and spin-readout, bringing us closer to realizing electron qubits in Paul traps.

### v2

# Image Current Detection of Electrons in a Room-Temperature Paul Trap

@@ Abstract
We report the image current detection of electrons in a room-temperature Paul trap at microwave frequencies. By selectively leveraging distinct cavity modes for trapping and detection, our approach efficiently extracts electron signals otherwise buried in the microwave drive used for pseudo-potential formation. When the trapped electrons resonate with the cavity mode, we observe a mode excitation and its exponential decay attributed to resistive cooling. Detuning electrons from the cavity resonance halts this decay, and sweeping electrons’ secular frequency reveals their oscillatory spectrum. Implementing this experiment at cryogenic temperatures could enable the image current detection and ground-state cooling of a single electron in Paul traps.
@@ I Introduction
Various trapping mechanisms have been explored to implement electron-based qubits in Penning traps. Planar Penning traps have gained attention due to their potential for two-dimensional scalability [26, 27, 28], which is required for quantum error correction algorithms such as surface codes [29]. However, their inherent lack of symmetry introduces significant anharmonicity in the trap potential, making image current detection of a single electron infeasible [30]. While theoretical methods have been proposed to mitigate this issue [31], no successful single electron measurements have been reported.
@@
Unlike Penning traps, where an electron exhibits large-radius magnetron motion, Paul traps confine an electron closer to the trap center where anharmonicity is minimal. This localization enhances the feasibility of both single electron detection and scalable two-dimensional architectures, shifting the attention toward trapping electrons in Paul traps [32, 33, 34, 35, 36]. This was recently demonstrated at room temperature [37], with detection achieved by ejecting trapped electrons onto a micro-channel plate.
@@
Essentially, quantum computation demands repeated operations, requiring the continuous confinement of electrons and their image current detection [5]. Equally crucial is ground-state cooling of the electron’s motion, a prerequisite for precise quantum manipulations [38]. However, both approaches present challenges unique to Paul traps: the former requires stringent filtering of the trap’s microwave field to extract the electron signal and protect detection circuits [35], while the latter demands resistive cooling at cryogenic temperatures, necessitating Paul traps inside a cryostat where cooling capacity is limited [37]. Overcoming these obstacles requires an integrated experimental framework combining advanced microwave filtering with power-efficient trap designs.
@@
Figure 1: Schematic illustration and characterization of the three-dimensional coupled coaxial cavities. (a) Design of the three-dimensional coupled coaxial cavities. (b) Cross-sectional view of the coupled coaxial cavities. The inset provides a magnified view of the pin geometry, where electrons are confined. (c) Radial ($r$-axis) electric field distribution of the out-of-phase $\lambda/4$ mode (readout mode) in cylindrical coordinates, used for image current detection of trapped electrons. The inset highlights the norm of electric fields between the pins, forming a localized electric dipole field. (d) Radial electric field distribution of the in-phase $45\lambda/$ 4 mode (trapping mode), which generates the pseudo-potential. The inset illustrates the corresponding potential structure responsible for electron confinement. (e) Top: Pseudo-potential $U_{\mathrm{eff}}$ along the $z$-axis, centered at potential minimum (blue curve), compared to an ideal harmonic potential (dashed orange curve). Bottom: Deviation $\Delta U_{\mathrm{eff}}$ from the ideal harmonic potential, quantifying the anahronicity. The gray regions indicate the positions of the pin tips. (f) Pseudo-potential distribution and its deviations along the radial axis which reveal the influence of the cavity geometry on the potential shape.
@@
To address these challenges, we previously proposed three-dimensional coupled coaxial cavities [38]. This device can efficiently extract electron signals from the trap’s microwave field by leveraging distinct resonance modes for pseudo-potential formation and image current detection. Moreover, this cavity concentrates electric fields at a single point and has a very high $Q$ factor when integrated with superconducting materials. Consequently, their synergy facilitates power-efficient electron Paul traps ($\sim\mathrm{mW}$) compatible with cryogenic operations. In this work, we successfully realize electron Paul traps using the coupled coaxial cavities. We achieve image current detection of electrons and observe exponential signal decay attributed to resistive cooling at room temperature. These results pave the way towards realizing image current detection and ground-state cooling of a single electron qubit at cryogenic temperature.
@@ II.1 Three-dimensional Coupled Coaxial Cavities
The pseudo-potential for the electron Paul trap is generated by oscillating quadrupole electric potential and can be formulated as $U_{\mathrm{eff}}=(q_{e}^{2}E^{2}/4m_{e}\Omega^{2})$. Here, $E$ and $\Omega$ represent the amplitude and angular frequency of the oscillating inhomogeneous electric field, while $q_{e}$ and $m_{e}$ are the charge and mass of the electron. Figure 1(e) and 1(f) show simulated pseudo-potentials along r- and z-axes when the pin voltages are $92\,\mathrm{V}$, which sets the electron oscillation to resonate with readout mode ($\omega_{z}=\omega_{1}^{\mathrm{out}}$). This pin voltage can be achieved with a microwave input power of $1.5\,\mathrm{W}$ under the current coupling condition and could be reduced to $800\,\mathrm{mW}$ if the critical coupling is achieved. The potential depth is approximately $0.8\,\mathrm{eV}$, limited by the confinement along the $z$-direction. The anharmonicity of the trap potential is illustrated as $\Delta_{\mathrm{eff}}$, the deviations from the ideal harmonic potentials (shown as orange dashed lines) in the bottom panels highlighted as green curves.
@@
We note that anharmonicity in the Pseudo-potential causes amplitude-dependent frequency shifts in the secular motion of trapped electrons. This effect broadens the motion spectrum and degrades the signal-to-noise ratio (SNR) in image current detection [30]. In our coupled coaxial cavities, anharmonicity primarily depends on the separation and the radius of the pin electrodes. For our experiment, we used a pin separation of $400\,\mathrm{\mu m}$ and a radius of $250\,\mathrm{\mu m}$. Considering the symmetry of the system, the pseudo-potential along the $z$-axis can be modeled using $U_{\mathrm{eff}}(z)=\frac{1}{2}m_{e}\omega_{z}^{2}(z^{2}+C_{4}z^{4}+C_{6}z^{6})$ ignoring odd-order terms. From the polynomial fit to the potential shown in Fig. 1(e), we obtained $C_{4}=-1.5\crossproduct 10^{-5}\,\mathrm{\mu m}^{-2}$ and $C_{6}\approx 0\,\mathrm{\mu m}^{-4}$. Therefore, the amplitude-dependent secular frequency shifts are given by [31]:
@@ II.2 Electron Source
To efficiently load electrons into traps, the kinetic energy of the loaded electrons must be sufficiently lower than the trap depth, highlighting the importance of preparing a low-energy electron source. In the Penning trap system, thermionic or field emission points are frequently used for this purpose [41, 42]. In these methods, high-energy primary electrons are generated first and collide with the device wall, producing background gases, and those background gases collide with other high-energy primary electrons, producing lower-energy secondary electrons that are supposed to be trapped.
This study used the photoelectron effect to generate high-energy primary electrons. We directed a continuous multimode laser with a central wavelength of $405\,\mathrm{nm}$ and a power of $300\,\mathrm{mW}$ onto the pins inside the coupled coaxial cavities through a hole [see Fig. 2(a)]. Although the energy of the laser at $405\,\mathrm{nm}$ (approximately $3.06\,\mathrm{eV}$) is theoretically insufficient to produce photoelectrons given the work function of gold plated on the device wall ($5.30\,\mathrm{eV}$) [43], biasing the pins with an oscillating electric field enables Schottky-enabled photoemission [44]. In addition, the multiple-photon process also facilitates photoemission [45].
@@ II.3 Experimental Setup
The signal of trapped electrons was observed by measuring the excitation of the readout mode via coupling ports $P_{2}$ and $P_{3}$ using a spectrum analyzer. Here, filtering the trap MW field from the electron signals is crucial to protect the amplifier and the spectrum analyzer. Therefore, we implemented a three-stage filtering in the electronic circuit. Firstly, coupling ports $P_{2}$ and $P_{3}$ are positioned at the node of the trapping mode, resulting in a $30\,\mathrm{dB}$ reduction of trap MW leakage into the detection circuit [Fig. 1(d)]. Secondly, coupling ports $P_{2}$ and $P_{3}$ are connected to the Double Arrow $180^{\circ}$ Hybrid Coupler to induce destructive interference of the trap MW [Fig. 2(a)]. While the trapping mode excites these ports in phase, the readout mode excites them with a phase difference of $\pi$. Since the hybrid coupler processes the input signals by outputting their sum or difference, an appropriate choice of output ports can suppress the signal from the trapping mode by $16\,\mathrm{dB}$ while preserving the electron signals. Finally, the signal is transmitted through two low-pass filters with a cut-off frequency of $1.65\,\mathrm{GHz}$, eliminating the trap MW by $80\,\mathrm{dB}$ [Fig. 2(c)]. These result in a total filtering of approximately $126\,\mathrm{dB}$ for trap MW. After the filtering, the readout mode excitation signals are amplified by $62\,\mathrm{dB}$ with a low-noise amplifier followed by another amplifier before the signal analyzer [Fig. 2(c)].
@@ III.1 Image Current Detection
When the $V_{\mathrm{MW}}$ was appropriately tuned to resonate electrons’ center-of-mass (COM) motion with the readout mode, an excitation of the readout mode was observed. Figure 3(b) presents the results for sequence (i), where $V_{\mathrm{MW}}$ remains constant to keep resonating electrons with the readout mode. The excitation signal appeared and gradually decayed with a time constant of $t_{\mathrm{cool}}=1.74\ \mathrm{s}$ before reaching a steady state. The observed excitation lasted over ten seconds even after electron loading finished, confirming the successful Paul trap of electrons and their image current detection.
@@ IV Conclusions and Outlook
The next challenge toward realizing electron qubits in Paul traps is the image current detection and spin-readout of a single electron at cryogenic temperature [34]. However, the anharmonicity in the trap potential remains a major obstacle. In Penning traps, this issue is mitigated through tailored static electric field compensation [1]. In contrast, applying such compensation in Paul traps has been challenging due to the complexity of the nonlinear Mathieu equation and limited to systems with small stability parameters ($q\ll 0.1$) [48]. Recently, we proposed a novel approach that extends static electric field compensation to nonlinear Mathieu equations, enabling anharmonicity suppression of Paul traps even in high stability parameters ($q>0.5$) [49]. Implementing this method could improve electron confinement and facilitate single-electron detection and spin-readout, bringing us closer to realizing electron qubits in Paul traps.